OpenStudy (anonymous):

Solve x/x-2 + 1/x-4 = 2/x^2-6x+8

OpenStudy (anonymous):

\[\frac{x}{x-2}+\frac{1}{x-4}=\frac{2}{x^2-6x+8}\]?

OpenStudy (anonymous):

the common denominator on the left is in fact \[x^2-6x+8\] what you get when you multiply the two denominators

OpenStudy (anonymous):

so you can solve \[x(x-4)+(x-2)=2\] for \(x\)